# Jinhang Chai     Jianqing Fan

### v1

@@ Appendix A Main Proofs
As $\bar{\mu}\geq\mu,\ \bar{r}\geq r,\ \bar{s}\geq s$, the feasibility of the ground-truth $P^{\star}$ follows from the form of optimization problem 3. Using the optimality condition, we have the following,
$\frac{1}{2}\|Y-\mathfrak{X}(\widehat{L}+\widehat{S})\|_{F}^{2}\leq\frac{1}{2}\|Y-\mathfrak{X}(L^{\star}+S^{\star})\|_{F}^{2}$ (20)
@@
$=\langle\mathfrak{X}^{\star}(W),\Delta_{L}+\Delta_{S}\rangle$
$=\langle\mathfrak{X}^{\star}(W),\Delta_{L}\rangle+\langle\mathfrak{X}^{\star}(W),\Delta_{S}\rangle$
$\leq\|\mathfrak{X}^{\star}(W)\|\cdot\|\Delta_{L}\|_{*}+\|\mathfrak{X}^{\star}(W)\|_{\max}\|\Delta_{S}\|_{1}$ (21)
where the first equality uses the definition of the conjugate operator and the last inequality uses Cauchy’s inequality.
Notice by the Assumption (3.4) and the definition of the feasible set of optimization problem (3), $\Delta_{L}$ has rank at most $r+\bar{r}$, $\|\Delta_{S}\|_{0}\leq s+\bar{s}$.
Therefore we have
$\|\Delta_{L}\|_{*}\leq\sqrt{r+\bar{r}}\|\Delta_{L}\|_{F}\leq\sqrt{2\bar{r}}\|\Delta_{L}\|_{F},\quad\|\Delta_{S}\|_{1}\leq\sqrt{s+\bar{s}}\|\Delta_{S}\|_{F}\leq\sqrt{2\bar{s}}\|\Delta_{S}\|_{F}$
Plug in (21), we get
$\frac{1}{2}\|\mathfrak{X}(\Delta_{L}+\Delta_{S})\|_{F}^{2}$ $\leq\|\mathfrak{X}^{\star}(W)\|\sqrt{2\bar{r}}\|\Delta_{L}\|_{F}+\|\mathfrak{X}^{\star}(W)\|_{\max}\sqrt{2\bar{s}}\|\Delta_{S}\|_{F}$
$\leq\sqrt{\|\Delta_{L}\|_{F}^{2}+\|\Delta_{S}\|_{F}^{2}}\sqrt{2\bar{r}\|\mathfrak{X}^{\star}(W)\|^{2}+2\bar{s}\|\mathfrak{X}^{\star}(W)\|_{\max}^{2}}$ (22)
@@
$\geq\frac{\kappa}{2}\|\Delta_{L}+\Delta_{S}\|_{\mathrm{F}}^{2}-2\tau(2\bar{r}\|\Delta_{L}\|_{F}^{2}+2\bar{s}\lambda^{2}\|\Delta_{S}\|_{F}^{2})$ (23)
@@
Now notice that $\Delta_{L}=\widehat{L}-L^{\star}=\widehat{U}\widehat{\Sigma}\widehat{V}^{\top}-U^{\star}\Sigma^{\star}V^{*T}$, where $\widehat{U},\widehat{V}\in\mathcal{O}_{p,\bar{r}}^{\bar{\mu}}$ and $U^{\star},V^{\star}\in\mathcal{O}_{p,r}^{\mu}\subset\mathcal{O}_{p,\bar{r}}^{\bar{\mu}}$.
By Lemma B.1, we have that
$\frac{\|\Delta_{L}\|^{2}_{\max}}{\|\Delta_{L}\|_{F}^{2}}\leq\frac{c\bar{\mu}\bar{r}^{4}}{p}$
Let $\operatorname{Supp}_{L}$ and $\operatorname{Supp}_{S}$ be the support of $\Delta_{L}$ and $\Delta_{S}$ respectively, denote $\mathcal{J}:=\operatorname{Supp}_{L}\bigcap\operatorname{Supp}_{S}$ and $\mathcal{T}_{\mathcal{J}}$ as the projection operator that zeros out the elements in $\mathcal{J}^{c}$. By definition we have $|\mathcal{J}|\leq 2\bar{s}\leq\frac{p}{4c\bar{\mu}\bar{r}^{4}}$.
Therefore, we can estimate the proportion of elements of $\Delta_{L}$ in $\mathcal{J}$ as follows,
$\frac{\|\mathcal{T}_{\mathcal{J}}(\Delta_{L})\|_{F}^{2}}{\|\Delta_{L}\|_{F}^{2}}\leq|\mathcal{J}|\frac{\|\Delta_{L}\|^{2}_{\max}}{\|\Delta_{L}\|_{F}^{2}}\leq\frac{c\bar{\mu}\bar{r}^{4}}{p}\frac{p}{4c\bar{\mu}\bar{r}^{4}}=\frac{1}{4}$
We have
@@
$\leq\frac{1}{4}\|\Delta_{L}\|_{F}^{2}+\frac{1}{4}\|\Delta_{S}\|_{F}^{2}$
@@
$\|\Delta_{L}+\Delta_{S}\|_{\mathrm{F}}^{2}=\|\Delta_{L}\|_{F}^{2}+\|\Delta_{S}\|_{F}^{2}+2\langle\Delta_{L},\Delta_{S}\rangle\geq\frac{1}{2}(\|\Delta_{L}\|_{F}^{2}+\|\Delta_{S}\|_{F}^{2})$
@@
$\geq\frac{\kappa}{8}(\|\Delta_{L}\|_{F}^{2}+\|\Delta_{S}\|_{F}^{2})$ (24)
Combining (A.1) and (A.1), we arrive at
$\|\Delta_{L}\|_{F}^{2}+\|\Delta_{S}\|_{F}^{2}\leq\frac{128}{\kappa^{2}}\left(\bar{r}\|\mathfrak{X}^{\star}(W)\|^{2}+\bar{s}\|\mathfrak{X}^{\star}(W)\|_{\max}^{2}\right),$ (25)
thus completes the proof. ∎
A.2 Proof of Theorem 3.7
@@
For the upper bound, we apply the deterministic result of our estimator for rectangular matrices which we denote here as $\mathcal{E}(Y)=(\widehat{L}^{\mathcal{E}},\widehat{S}^{\mathcal{E}})$. In the first noise scenario, we have by the condition on the sparsity level, the first term in the bound (5) dominates the second one and therefore, up to a constant,
$\mathfrak{R}^{1}_{\rho}\leq\mathbb{E}\rho(L,S,\widehat{L}^{\mathcal{E}},\widehat{S}^{\mathcal{E}})\lesssim\mathbb{E}r\|W\|^{2}\lesssim\max\{p,q\}r\sigma^{2}$
For the second case, the second term in the bound (5) dominates the first one and therefore, up to logarithmic factors,
@@
where the last inequality follows from a union bound.
Therefore, we are left with the proof of the lower bounds. The proof is based on the combination of Fano’s method along with the construction of local packing using Gilbert-Varshamov bound.
@@
We invoke the following Gilbert-Varshamov bound in Lemma E.1 which is included in Section E for completeness (for a proof of this lemma, see Yu, (1997)). Therefore, we get a subset $A_{1}$ of $\{-\eta,\eta\}^{p,r}$ (flattened to vectors), such that $|A_{1}|\geq\exp(c_{0}pr)$ and
@@
$\|(f(R_{1})-f(R_{2}))\|\\ _{F}^{2}=l_{1}\|R_{1}-R_{2}\|^{2}_{F}\geq l_{1}\eta^{2}\frac{4pr}{3}\geq\eta^{2}\frac{2pq}{3}$
On the other hand, the KL-divergence between the isotropic Gaussian distributions with mean $(f(R_{1})$ and mean $f(R_{2}))$ is upper bounded by $pq\frac{4\eta^{2}}{\sigma^{2}}$ where we use the chain rule for KL-divergence and KL-divergence for two univariate Gaussian random variables.
@@
$\geq\eta^{2}\frac{2pq}{3}\left(1-\frac{pq\frac{4\eta^{2}}{\sigma^{2}}+\log 2}{c_{0}pr}\right)$
@@
$\mathfrak{R}^{d_{1}}_{\rho}\geq\eta^{2}\frac{2pq}{9}\gtrsim pr\sigma^{2}=\max\{p,q\}r\sigma^{2}$
@@
$\gtrsim s\sigma^{2}$
A.3 Proof of Proposition 4.3
@@
To prove this proposition, we will need the following technical lemma whose proof is relegated to Section E.
Lemma A.1.
Let $Y=\{Y_{0},Y_{1},\cdots,Y_{n-1}\}$ be the new Markov chain defined in (27). Denote $\epsilon$-mixing time of the original Markov chain $X=\{X_{0},X_{1},\cdots,X_{n-1}\}$ as $\tau_{X}(\epsilon)$ and transition matrix as $P$. Similarly, denote $\epsilon$-mixing time of $Y$ as $\tau_{Y}(\epsilon)$ and transition matrix as $Q$. Suppose $X$ is ergodic, then $Y$ is ergodic and $\tau_{Y}(\epsilon)=\tau_{X}(\epsilon)$.
@@
Consider a new Markov chain that takes the consecutive states in the original chain as the new state. Formally, define $Y=\{Y_{0},Y_{1},\cdots,Y_{n-1}\}$ as the new Markov chain taking values in the product space $[p]\times[p]\backslash Z$, with $Q$ its transition matrix, where
@@
and $Z=\{(i,j)\big{|}P_{i,j}=0\}$
Now by Lemma A.1, $Y$ is ergodic. Also by (63),
@@
Applying (74) of the lemma 7 of Zhang & Wang, (2019) on the new Markov chain $Y$, we have for all $n\geq C\tau_{*}p\log^{2}n$,
@@
By the Lemma A.1, we have $\tau_{Y}(\frac{1}{4})=\tau_{X}(\frac{1}{4})=\tau_{*}$, and (10) holds, thus completing the proof. ∎
A.4 Proof of Theorem 4.4
@@
First of all, notice that it suffices to prove (11). This is because (12) follows from $\|\widehat{F}-F^{\star}\|_{1}\leq p\|\widehat{F}-F^{\star}\|_{F}$, (14) follows from $\|\widehat{P}-P^{\star}\|_{1}\leq\frac{\|\widehat{F}-F^{\star}\|_{1}}{\pi_{\min}}$, and (13) follows from $\|\widehat{P}-P^{\star}\|_{F}\leq\frac{\|\widehat{F}-F^{\star}\|_{F}}{\pi_{\min}}$.
@@
$\leq c_{v}\operatorname{Tr}\left((\widehat{P}-P)^{\top}(\widehat{P}-P)\right)=c_{v}\|\widehat{P}-P\|_{F}^{2}$
@@
as desired. Further, if $p_{\max}=\operatorname{O}(\frac{1}{p})$, $\pi_{\max}\asymp\pi_{\min}$ and $\tau_{*}=\operatorname{O}(1)$, we have $c_{v}\frac{c\pi_{\max}\tau_{*}\log^{2}n}{n\pi_{\min}^{2}}(r+p_{\max}s)=\operatorname{O}\left(c_{v}\frac{rp}{n}\right)$. ∎
A.6 Proof of Theorem 4.7
@@
Appendix B Key Separation Lemma
In this section, we aim to prove the key separation Lemma B.1 which claims the difference between two incoherent low-rank matrices should distribute relatively evenly across elements. In other words, incoherent low-rank matrices are in some sense well-separated. Section B.1 lists a number of intermediate lemmas to achieve Lemma B.1. Section B.6 is devoted to the proof of Lemma B.1. In Section B.8, we discuss the optimality of Lemma B.1 with respect to the order of dimension $p$. With slight abuse of notation, the $L$ in this section is referred to as the left orthogonal part of $Q^{\prime}s$ SVD instead of the low-rank part in the main text. To prevent confusion, all notations used in this section do not apply to other sections except D.
Lemma B.1.
Let $P=U\Sigma V^{\top}$ and $Q=L\Lambda R^{\top}$ be the SVD of $P,Q\in\mathbb{R}^{p\times p}$, $U,V,L,R\in\mathcal{O}_{p,r}$. Denote $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$ and $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{r}\geq 0$ as the diagonal elements of $\Sigma$ and $\Lambda$, respectively. Suppose $P$ and $Q$ both satisfy the $\mu$-incoherence condition 3.3, then we have for $\Delta=P-Q$,
$\frac{\|\Delta\|_{\max}^{2}}{\|\Delta\|_{F}^{2}}\leq\frac{\tilde{c}\mu r^{4}}{p}$ (28)

### v2

@@ Appendix B Main Proofs in Section 2
The proof highly relies on the key separation lemma D.1 whose detailed proof is given in Appendix D.
As $\bar{\mu}\geq\mu,\ \bar{r}\geq r,\ \bar{s}\geq s$, the feasibility of the ground-truth $P^{\star}$ follows from the form of optimization problem (3). Using the optimality condition, we have the following,
$\frac{1}{2}\|Y-\mathfrak{X}(\widehat{L}+\widehat{S})\|_{F}^{2}\leq\frac{1}{2}\|Y-\mathfrak{X}(L^{\star}+S^{\star})\|_{F}^{2}.$ (24)
@@
$=\langle\mathfrak{X}^{*}(W),\Delta_{L}+\Delta_{S}\rangle$
$=\langle\mathfrak{X}^{*}(W),\Delta_{L}\rangle+\langle\mathfrak{X}^{*}(W),\Delta_{S}\rangle$
$\leq\|\mathfrak{X}^{*}(W)\|\cdot\|\Delta_{L}\|_{*}+\|\mathfrak{X}^{*}(W)\|_{\max}\|\Delta_{S}\|_{1},$ (25)
where the first equality uses the definition of the conjugate operator, and the last inequality uses the Cauchy-Schwarz inequality.
Notice by the Assumption (2.4) and the definition of the feasible set of optimization problem (3), $\Delta_{L}$ has rank at most $r+\bar{r}$, $\|\Delta_{S}\|_{0}\leq s+\bar{s}$. Therefore we have
$\|\Delta_{L}\|_{*}\leq\sqrt{r+\bar{r}}\|\Delta_{L}\|_{F}\leq\sqrt{2\bar{r}}\|\Delta_{L}\|_{F},\quad\|\Delta_{S}\|_{1}\leq\sqrt{s+\bar{s}}\|\Delta_{S}\|_{F}\leq\sqrt{2\bar{s}}\|\Delta_{S}\|_{F}.$
Plugging in (25), we get
$\frac{1}{2}\|\mathfrak{X}(\Delta_{L}+\Delta_{S})\|_{F}^{2}$ $\leq\|\mathfrak{X}^{*}(W)\|\sqrt{2\bar{r}}\|\Delta_{L}\|_{F}+\|\mathfrak{X}^{*}(W)\|_{\max}\sqrt{2\bar{s}}\|\Delta_{S}\|_{F}$
$\leq\sqrt{\|\Delta_{L}\|_{F}^{2}+\|\Delta_{S}\|_{F}^{2}}\sqrt{2\bar{r}\|\mathfrak{X}^{*}(W)\|^{2}+2\bar{s}\|\mathfrak{X}^{*}(W)\|_{\max}^{2}}.$ (26)
@@
$\geq\frac{\kappa}{2}\|\Delta_{L}+\Delta_{S}\|_{\mathrm{F}}^{2}-2\tau(2\bar{r}\|\Delta_{L}\|_{F}^{2}+2\bar{s}\lambda^{2}\|\Delta_{S}\|_{F}^{2}),$ (27)
@@
Now notice that $\Delta_{L}=\widehat{L}-L^{\star}=\widehat{U}\widehat{\Sigma}\widehat{V}^{\top}-U^{\star}\Sigma^{\star}V^{*T}$, where $\widehat{U},\widehat{V}\in\mathcal{O}_{p,\bar{r}}^{\bar{\mu}}$ and $U^{\star},V^{\star}\in\mathcal{O}_{p,r}^{\mu}\subset\mathcal{O}_{p,\bar{r}}^{\bar{\mu}}$. By Lemma D.1, we have that
$\frac{\|\Delta_{L}\|^{2}_{\max}}{\|\Delta_{L}\|_{F}^{2}}\leq\frac{c\bar{\mu}\bar{r}^{3}}{p}.$
Let $\operatorname{Supp}_{L}$ and $\operatorname{Supp}_{S}$ be the support of $\Delta_{L}$ and $\Delta_{S}$ respectively, denote $\mathcal{J}:=\operatorname{Supp}_{L}\bigcap\operatorname{Supp}_{S}$ and $\mathcal{T}_{\mathcal{J}}$ as the projection operator that zeros out the elements in $\mathcal{J}^{c}$. By the definition of ${\cal J}$ and Assumption 2.4 we have $|\mathcal{J}|\leq 2\bar{s}\leq\frac{p}{4c\bar{\mu}\bar{r}^{3}}$ where we take $\bar{c}=8c$. Therefore, we can estimate the proportion of elements of $\Delta_{L}$ in $\mathcal{J}$ as follows,
$\frac{\|\mathcal{T}_{\mathcal{J}}(\Delta_{L})\|_{F}^{2}}{\|\Delta_{L}\|_{F}^{2}}\leq|\mathcal{J}|\frac{\|\Delta_{L}\|^{2}_{\max}}{\|\Delta_{L}\|_{F}^{2}}\leq\frac{c\bar{\mu}\bar{r}^{3}}{p}\frac{p}{4c\bar{\mu}\bar{r}^{3}}=\frac{1}{4}.$
As a result, we have
@@
$\leq\frac{1}{4}\|\Delta_{L}\|_{F}^{2}+\frac{1}{4}\|\Delta_{S}\|_{F}^{2},$
@@
$\|\Delta_{L}+\Delta_{S}\|_{\mathrm{F}}^{2}=\|\Delta_{L}\|_{F}^{2}+\|\Delta_{S}\|_{F}^{2}+2\langle\Delta_{L},\Delta_{S}\rangle\geq\frac{1}{2}(\|\Delta_{L}\|_{F}^{2}+\|\Delta_{S}\|_{F}^{2}).$
@@
$\geq\frac{\kappa}{8}(\|\Delta_{L}\|_{F}^{2}+\|\Delta_{S}\|_{F}^{2}).$ (28)
Combining (B.1) and (B.1), we arrive at
$\|\Delta_{L}\|_{F}^{2}+\|\Delta_{S}\|_{F}^{2}\leq\frac{128}{\kappa^{2}}\left(\bar{r}\|\mathfrak{X}^{*}(W)\|^{2}+\bar{s}\|\mathfrak{X}^{*}(W)\|_{\max}^{2}\right),$ (29)
thus completing the proof. ∎
B.2 Proof of Theorem 2.8
@@
Upper Bound: For the upper bound, we apply the deterministic result of our estimator for rectangular matrices, which we denote here as $\mathcal{E}(Y)=(\widehat{L}^{\mathcal{E}},\widehat{S}^{\mathcal{E}})$. In the first noise scenario, we have by the condition on the sparsity level, the first term in the bound (5) dominates the second one and therefore, up to a constant,
$\mathfrak{R}^{1}_{\rho}\leq\mathbb{E}\rho(L,S,\widehat{L}^{\mathcal{E}},\widehat{S}^{\mathcal{E}})\lesssim\mathbb{E}r\|W\|^{2}_{2}\lesssim\max\{p,q\}r\sigma^{2},$
where the last inequality bounds $\mathbb{E}\|W\|_{2}$ and follows in the same vein from Theorem 5.39 of Vershynin, (2010), except now we have different covariance matrices across rows.
For the second case, denote the nonzero index set of $E$ as ${\cal J}_{E}$. For $t>0$, we have by union bound that with probability at least $1-2se^{-t^{2}/2}$ that $|W_{i,j}|\leq\sigma t$ for every $(i,j)\in{\cal J}_{E}$. We denote this event as ${\cal A}$. It holds that under ${\cal A}$, $\|W\|\leq\sigma t\|E\|$. Hence we have
$\mathbb{E}[\|W\|^{2}]$ $=\mathbb{E}[\|W\|^{2}{\cal I}_{\cal A}]+\mathbb{E}[\|W\|^{2}{\cal I}_{{\cal A}^{c}}].$
Moreover, we have the bound $\mathbb{E}[\|W\|^{2}{\cal I}_{\cal A}]\leq\sigma^{2}t^{2}\|E\|^{2}$, as well as
$\mathbb{E}[\|W\|^{2}{\cal I}_{{\cal A}^{c}}]$ $\leq\mathbb{E}[\|W\|_{F}^{2}{\cal I}_{{\cal A}^{c}}]$
$\leq\sum_{(i,j)\in{\cal J}_{E}}\left\{\mathbb{E}[W_{i,j}^{2}{\cal I}(|W_{i,j}|\geq\sigma t)]+\sum_{(i^{\prime},j^{\prime})\in{\cal J}_{E}\backslash(i,j)}\mathbb{E}[W_{i,j}^{2}{\cal I}(|W_{i^{\prime},j^{\prime}}|\geq\sigma t)]\right\}$
$\lesssim s^{2}\sigma^{2}(t+1)e^{-t^{2}/2},$
where we used that $\int_{t}^{\infty}x^{2}e^{-x^{2}/2}\lesssim(t+1)e^{-t^{2}/2}$. Take $t=C\log(s)$ with sufficiently large constant $C$, we have $\mathbb{E}[\|W\|^{2}]\lesssim\sigma^{2}\|E\|^{2}$ up to log terms. And hence $r\mathbb{E}[\|W\|^{2}]\lesssim s\sigma^{2}\leq s\mathbb{E}[\|W\|_{\max}^{2}]$ by the assumption on $\|E\|$.
Therefore, the second term in the bound (5) dominates the first one and, therefore, up to logarithmic factors,
@@
where the bound on $\|W\|_{\max}$ follows from a similar union bound.
Lower Bound: We are left with the proof of the lower bounds. The proof is based on the combination of Fano’s method along with the construction of local packing using Gilbert-Varshamov bound.
@@
We invoke the following Gilbert-Varshamov bound in Lemma F.1 in the supplementary material which is included in Appendix F in the supplementary material for completeness (for a proof of this lemma, see Yu, (1997)). Therefore, we get a subset $A_{1}$ of $\{-\eta,\eta\}^{p,r}$ (flattened to vectors), such that $|A_{1}|\geq\exp(c_{0}pr)$ and
@@
$\|(f(R_{1})-f(R_{2}))\|\\ _{F}^{2}=l_{1}\|R_{1}-R_{2}\|^{2}_{F}\geq l_{1}\eta^{2}\frac{4pr}{3}\geq\eta^{2}\frac{2pq}{3}.$
On the other hand, the KL-divergence between the isotropic Gaussian distributions with mean $f(R_{1})$ and mean $f(R_{2})$ is upper bounded by $\sum_{i=1}^{p}[f(R_{1})-f(R_{2})]_{i}^{T}\Sigma_{i}^{-1}[f(R_{1})-f(R_{2})]_{i}\lesssim\frac{pq\eta^{2}}{\sigma^{2}}$, where we use the chain rule for KL-divergence and KL-divergence for two multivariate Gaussian random variables with same covariance $\Sigma$ is given by $KL({\cal N}(\mu+\delta,\Sigma)|{\cal N}(\mu,\Sigma))=\mu^{T}\Sigma^{-1}\mu/2$.
@@
$\geq\eta^{2}\frac{2pq}{3}\left(1-\frac{pq\frac{4\eta^{2}}{\sigma^{2}}+\log 2}{c_{0}pr}\right).$
@@
$\mathfrak{R}^{d_{1}}_{\rho}\geq\eta^{2}\frac{2pq}{9}\gtrsim pr\sigma^{2}=\max\{p,q\}r\sigma^{2}.$
@@
$\gtrsim s\sigma^{2}.$
Appendix C Proofs in Section 3
C.1 Proof of Proposition 3.2
@@ Appendix C Proofs in Section 3
To prove the second part, consider a new Markov chain that takes the consecutive states in the original chain as the new state. Formally, define $Y=\{Y_{0},Y_{1},\cdots,Y_{n-1}\}$ as the new Markov chain taking values in the product space $[p]\times[p]\backslash Z$, with $Q$ its transition matrix, where
@@
and $Z=\{(i,j)\big{|}P_{i,j}=0\}$. To continue, we will need the following technical lemma whose proof is relegated to Section G.
Lemma C.1.
Let $Y=\{Y_{0},Y_{1},\cdots,Y_{n-1}\}$ be the new Markov chain defined in (31). Denote the $\epsilon$-mixing time of the original Markov chain $X=\{X_{0},X_{1},\cdots,X_{n-1}\}$ as $\tau_{X}(\epsilon)$ and transition matrix as $P$. Similarly, denote $\epsilon$-mixing time of $Y$ as $\tau_{Y}(\epsilon)$ and transition matrix as $Q$. Suppose $X$ is ergodic, then $Y$ is ergodic and $\tau_{Y}(\epsilon)=\tau_{X}(\epsilon)$.
Now by Lemma C.1, it holds that $Y$ is ergodic. Also by (62),
@@
Applying (74) of Lemma 7 of Zhang & Wang, (2019) on the new Markov chain $Y$, we have for all $n\geq C\tau_{*}p\log^{2}n$,
@@
By the Lemma C.1, we have $\tau_{Y}(\frac{1}{4})=\tau_{X}(\frac{1}{4})=\tau_{*}$, and (13) holds, thus completing the proof. ∎
C.2 Proof of Theorem 3.3
@@
First of all, notice that it suffices to prove (14). As a matter of fact, (15) follows from $\|\widehat{F}-F^{\star}\|_{1}\leq p\|\widehat{F}-F^{\star}\|_{F}$. And by applying Lemma 2 in Zhang & Wang, (2019) to the rows of $\widehat{P}-P^{\star}$ and summing up, we have $\|\widehat{P}-P^{\star}\|_{1}\leq\frac{\|\widehat{F}-F^{\star}\|_{1}}{\pi_{\min}}$ which implies (16).
@@
$\leq c_{v}\operatorname{Tr}\left(\mathbb{E}(\widehat{P}-P)^{\top}(\widehat{P}-P)\right)=c_{v}\mathbb{E}\|\widehat{P}-P\|_{F}^{2}$
@@
as desired. Further, if $p_{\max}=\operatorname{O}(\frac{1}{p})$, $\pi_{\max}\asymp\pi_{\min}$ and $\tau_{*}=\operatorname{O}(1)$, we have $c_{v}\frac{c\pi_{\max}\tau_{*}\log^{2}n}{n\pi_{\min}^{2}}(r+p_{\max}s)=\operatorname{O}\left(c_{v}\frac{rp\log^{2}n}{n}\right)$. ∎
C.4 Proof of Theorem 3.6
@@
Appendix D Key Separation Lemma
In this section, we aim to prove the key separation Lemma D.1 which claims the difference between two incoherent low-rank matrices should distribute relatively evenly across elements. In other words, incoherent low-rank matrices are in some sense well-separated. Section D.1 lists a number of intermediate lemmas to achieve Lemma D.1. Section D.6 is devoted to the proof of Lemma D.1. In Section D.8, we discuss the optimality of Lemma D.1 with respect to the order of dimension $p$. With slight abuse of notation, the $L$ in this section is referred to as the left orthogonal part of $Q^{\prime}s$ SVD instead of the low-rank part in the main text. To prevent confusion, all notations used in this section do not apply to other sections except Section F.
Lemma D.1.
Let $P=U\Sigma V^{\top}$ and $Q=L\Lambda R^{\top}$ be the SVD of $P,Q\in\mathbb{R}^{p\times p}$, $U,V,L,R\in\mathcal{O}_{p,r}$. Denote $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{r}\geq 0$ and $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{r}\geq 0$ as the diagonal elements of $\Sigma$ and $\Lambda$, respectively. Suppose $P$ and $Q$ both satisfy the $\mu$-incoherence condition 2.3, then we have for $\Delta=P-Q$,
$\frac{\|\Delta\|_{\max}^{2}}{\|\Delta\|_{F}^{2}}\leq\frac{\tilde{c}\mu r^{3}}{p}$ (32)